The Washington Post reported on 25thFeb 2019

that Los Angeles subway ridership declined after a fare increase: “There were nearly 4 million fewer riders in December 2018, the first full month after the price of a token increased 25 cents to $1.50 than in the previous December, a 4.3 percent decline.”


a. Use these data to estimate the price elasticity of demand for subway rides.

b. According to your estimate, what happens to the Transit Authority of LA’s revenue when the fare rises?

c. Why might your estimate of the elasticity be unreliable?




1
Expert's answer
2021-10-20T15:56:14-0400

(a)

The percentage change in quantity is 4.3%.

The percentage change in price: "=\frac{1.50-1.25}{1.25}\times100=20" %.

Price Elasticity of Demand:"= \frac{Percentage Change in Quantity}{Percentage Change in Price}"

"=\frac{4.3}{20}"

"=0.22."

(b)

Revenue will increase when fare rises because the demand is inelastic as indicated by the value of elasticity of demand above.

(c)

The estimate of elasticity might be unreliable because it is only the first month after the increase in fare. As time goes by, people may switch to other means of transport due to the increase in fare. So, in the Long run, the elasticity might be larger than in the short run.